OpenStudy (anonymous):

If secθ=5/3, find cotθ.

OpenStudy (anonymous):

\[\text{Cot}\left[\text{ArcSec}\left[\frac{5}{3}\right]\right]=\frac{3}{4} \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your welcome and than your for the medal.